how would you solve a system of equations algebracially that is 5x+2y=2 and x-7y=-81?
Answer by checkley77(12844)

5x+2y=2
x-7y=-81 MULTIPLY BY -5 & ADD.
-5X+35Y=405
5X+2Y=2
---------------
37Y=407
Y=407/37
Y=11 ANS.
5X+2*11=2
5X+22=2
5X=2-22
5X=-20
X=-20/5
X=-4 ANS.
PROOF:
-4-7*11=-81
-4-77=-81
-81=-81